# Coherent synchrotron radiation by excitation of surface plasmon polariton on near-critical solid microtube surface

### v1

@@ Abstract
Coherent synchrotron radiation (CSR) is vital for developing powerful ultrashort light sources. We introduce a CSR generation mechanism using surface plasmon polaritons (SPPs) resonantly excited on a solid, near-critical-density microtube. A high-intensity, circularly polarised laser pulse, propagating along the microtube axis, efficiently couples the cylindrical SPP modes. This process creates azimuthally structured, rotating electromagnetic fields. These rotating fields subsequently confine, modulate, and directly accelerate surface electrons, causing them to emit CSR in the Valilov-Cherenkov angle. We further demonstrate that by improving the azimuthal symmetry, the helical modulation enables CSR emission across all azimuthal directions, significantly enhancing radiation intensity even when full coherence is imperfect. The harmonics can be well isolated for a high charge beam. Our full 3D Particle-in-Cell simulations indicate this scheme can generate X-rays with coherence enhanced by up to two orders of magnitude compared to incoherent emission.
@@
Coherent synchrotron radiation (CSR) represents a unique regime of electromagnetic (EM) radiation emitted when relativistic charged particles—typically electrons—undergo accelerated motion in a collective, spatially ordered configuration [1]. Unlike conventional incoherent synchrotron radiation, where individual particles emit independently, CSR arises from the constructive interference of EM fields at specific wavelengths with fixed phase relationships between radiating electrons. The coherence results in a quadratic increase in power with the number of electrons, rather than the linear increase seen in incoherent cases. This property makes CSR highly advantageous for applications requiring high-power radiation sources in the terahertz (THz) and far-infrared regimes, such as ultrafast spectroscopy [2], particle acceleration diagnostics [3, 4], and next-generation light sources [5, 6].
Figure 1: Schematic: a CP laser pulse (yellow) enters the vacuum channel of the microtube (grey) and excites SPPs while being scattered at the sharp edge. The rotation mode $m=\pm 1$ (green-yellow) can efficiently couple with the laser field and accelerates the trapped electrons(sphere, blue-to-red colour presents the energy) into a spiral motion along the solid surface for the CSR generation (yellow-red cone).
@@
The rapid progress in relativistic laser-plasma interactions has promoted the emergence of new compact radiation sources [7, 8, 9, 10]. Furthermore, an unexplored regime of superradiance has been recently discovered where coherent radiation can be obtained by specific modulation of the bunch to make the quasiparticle travel luminally or superluminally [11], which is distinct from the conventional superradiance mechanism [12]. This superluminal motion causes radiated phase fronts emitted at different times to intersect, forming an optical shock or caustic. At the crossing angle, the spherical waves emitted by other particles in the bunch have the same phase and thus interfere constructively. This new theory unlocks novel CSR in plasma-based accelerators [13].
@@
A high-intensity laser pulse interacting with an interface between a metal and a dielectric medium (e.g., air or vacuum) can induce collective oscillations of free electrons. These oscillations, known as surface plasmons (SPs), are characterised by the plasma frequency, $\omega_{p}=\sqrt{4\pi n_{e}e^{2}/m_{e}}$, where $n_{e}$ is electron density, $e$ is electron charge, $m_{e}$ is electron rest mass [14]. The excitation of SPs has been recently proposed for generating TV/m-level wakefields in solid microtubes for relativistic particle acceleration [15]. When these SPs couple with incident EM waves, they form surface plasmon polaritons (SPPs)—hybrid light-electron modes that propagate along the interface, also termed surface plasmon waves [16]. Direct excitation of SPPs on a smooth, infinite interface is inherently inefficient due to the momentum mismatch between the incident light and the SPP modes [17]. However, in finite-sized solid microtargets, sharp edges can serve as effective scatterers, supplying the necessary momentum to bridge this gap and facilitate efficient SPP excitation [18]. The EM fields associated with SPPs can accelerate electrons along the target surface to relativistic energies. Recent studies have shown that irradiating the edge of a flat microtape with a high-intensity, linearly polarised laser pulse can generate strong SPP, producing high-energy electrons and high-flux X-ray radiation [19].
@@
In this letter, we first show that SPPs can be directly excited at the cylindrical solid interface between plasma and vacuum by a circularly polarised (CP) laser pulse propagating paraxially through a microtube of near-critical density, as illustrated in Fig. 1. The cylindrical geometry enables the CP laser to couple efficiently with a rotating SPP mode whose phase velocity approaches the speed of light in vacuum. This resonant coupling provides $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP fields for the modulated trapping, confinement and stable acceleration of relativistic electrons over several Rayleigh lengths. By matching the channel radius with the SPP field, the modulated high-energy electrons stably rotate near the surface and can produce CSR in the Valilov-Cherenkov angle. We further show that with an azimuthally symmetric distribution, the modulated beam can emit coherent radiation in all azimuthal directions. It also indicates that the harmonics can be well isolated. Full 3D particle-in-cell (PIC) simulations have been used to verify the fundamental process. Numerical simulations show significant coherence enhancement of photon emission in comparison with the incoherent case, with the radiation spectrum of critical energy at tens of $\mathrm{k}\mathrm{e}\mathrm{V}$.
A microtube can be made of nano-materials, e.g. longitudinally aligned carbon nanotube (CNT) forest, which can provide the solid plasma with flexible density in the range of $10^{19}-10^{24}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ [20]. The SPP field is shown in Fig. 2, which is obtained from a PIC simulation where a microtube of radius $a=1.5\leavevmode\nobreak\ $\mu\mathrm{m}$$ is used. A $20\leavevmode\nobreak\ $\mu\mathrm{m}$$-long vacuum section is placed at the head of the target to initialise the laser pulse and provides a sharp annular edge. As the bulk size of the tube is much larger than the inner radius of a CNT, e.g. up to several $\mathrm{n}\mathrm{m}$, the effect from the inner structure of CNTs is neglected. A CP laser pulse of wavelength $\lambda_{l}=0.8\leavevmode\nobreak\ $\mu\mathrm{m}$$ and normalized strength $a_{0}=6$ propagates in the $z$ direction and is focused to the entrance of the microtube at $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$, with root-mean-square (RMS) waist $w_{0}=2.0\leavevmode\nobreak\ $\mu\mathrm{m}$$, and duration $\sigma_{\tau}=20.0\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$. These parameters give a peak field of $24.0\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$, or $13.6\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ on the inner surface. This is sufficient for the pulse front to instantaneously ionise electrons of the material within half a cycle. Therefore, the material properties, e.g., lattice structure or optical absorption, are unimportant for electron dynamics. For ultrashort laser pulses, e.g. duration $L\ll 100\leavevmode\nobreak\ $\mathrm{f}\mathrm{s}$$, the plasma expansion can be neglected and the solid target can be assumed to maintain the sharp-edge cold plasma profile. These conditions allow the propagation of SPPs on the vacuum-plasma interface.
The simulations are carried out with the code WarpX [21]. The dimensions of the moving window are $8\leavevmode\nobreak\ $\mu\mathrm{m}$\times 8\leavevmode\nobreak\ $\mu\mathrm{m}$\times 30\leavevmode\nobreak\ $\mu\mathrm{m}$$, comprising $512\times 512\times 1024$ cells in the $x$, $y$, and $z$ directions, respectively. Each cell contains 64 macro particles, which are sufficient to solve the electron dynamics effectively. The microtube is initially modelled as a uniform distribution of neutral carbon atoms, with their ionisation energies adjusted to reflect CNT properties—specifically the weak $\pi$-bond and $\sigma$ (C–C) bond energies. The initial carbon density is $n_{C}=2\times 10^{21}\leavevmode\nobreak\ $\mathrm{c}\mathrm{m}^{-3}$$ which gives $\omega_{p}$ close to the laser frequency $\omega_{l}$. The field ionisation is implemented using the Ammosov-Delone-Krainov (ADK) method  [22].
@@
Figure 2: (a) Electron density distribution $n_{e}$ (black colour) and $E_{z}$ (blue-red colourmap) in $yz$-plane sliced at $y=0$ while the laser pulse entering the microtube, respectively. (b) Snapshots of $E_{z}$ in $xy$-plane sliced at $z=5\leavevmode\nobreak\ $\mu\mathrm{m}$$ corresponding to (a). (c)-(e) Snapshots of $E_{z}$, $E_{r}$ and $E_{\phi}$ in $xy$-plane at the same phase after the laser pulse propagating $z=41\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the tube. The solid black and dashed red lines in (e) present the line plots of $E_{z}$ along $y=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ from PIC and analytical solutions, respectively.
@@
The vertical solid surface at $z=0\leavevmode\nobreak\ $\mu\mathrm{m}$$ scatters the incident CP laser pulse and enables the direct matching with the SPP mode if the resonant condition $k_{l}=k_{sp}$ are satisfied, where $k_{l}$ and $k_{sp}$ are wave numbers of laser and SPP field, respectively. This condition is efficient near critical density, as shown in Fig. 2(a) and (b). The excited SPP fields can stably propagate with ultrahigh strength in $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level, e.g. $4.5\leavevmode\nobreak\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$$ in our simulations as shown in Fig. 2 (c)-(e) where the components are sliced after the laser pulse propagating $z=41\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the microtube. Some electrons can be trapped into the SPP field and accelerated to relativistic speed by the longitudinal component $E_{z}$ and, at the same time, confined in a small vicinity near the surface by the transverse component $E_{r}$ if they have sufficient energy, as shown in Fig. 2 (c). Due to the rotating phase front of the CP laser pulse, the electron beam is modulated to be helical during the trapping process, as shown in Fig. 3(a). The azimuthal component $E_{\phi}$ drives the modulated beam to rotate.
@@
The PIC results can be confirmed theoretically by analysing the SPP excitation in cylindrical geometry, $(r,\phi,z)$ [23]. In general, a pure transverse magnetic (TM) mode (e.g. $B_{spp,z}=0$ and $E_{spp,z}\neq 0$) of SPP field can exist on the cylindrical surface. The $z$-component of a SPP mode can be given by solving Maxwell equations, as $E_{spp,z}^{(m)}(r,\phi,z)=Af_{m}(\kappa r)e^{i(k_{z}z+m\phi-\omega t)}$, where $A$ is a constant representing the field amplitude. $m$ is the SPP mode number and $\kappa=\sqrt{k_{z}^{2}-\epsilon_{d}\omega^{2}/c^{2}}$ with $k_{z}$ the wavevector along the tube axis, $\omega$ the frequency, $c$ is the speed of light in vacuum and $\epsilon_{d}$ permittivity of the medium. $f_{m}(\kappa r)$ presents the radial field profile and depends on the medium properties. In the vacuum channel, $r<a$, the field is presented by the modified Bessel functions of the first kind as $f_{m}(\kappa r)=I_{m}(\kappa_{1}r)$ since the field is finite as $r\to 0$ and $\epsilon_{d}=\epsilon_{1}=1$. In the plasma, $r>a$, the field is presented by the the modified Bessel functions of the second kind as $f_{m}(\kappa r)=K_{m}(\kappa r)$ the field decays to zero as $r\to\infty$ and $\epsilon_{d}=\epsilon_{2}=1-n_{e}/n_{c}(\omega)$ where $n_{c}$ is the critical density for the frequency $\omega$ as $n_{c}(\omega)=m_{e}\omega^{2}/4\pi e^{2}$. PIC simulation agrees well with the analytical solution as shown in Fig. 2(c)-(e). The other components can be obtained from the axial components using Maxwell’s equations, for example, the electric components as $E_{ssp,r}^{(m)}=(ik_{z}/\kappa^{2})\partial E_{spp,z}^{(m)}/\partial r$ and $E_{spp,\phi}^{(m)}=imk_{z}E_{spp,z}^{(m)}/\kappa^{2}r$ for different regions, respectively. See more details in the Supplementary Material.
The dispersion relation can be derived by considering the continuity of the SPP field at the interface of plasma and vacuum as
where $\kappa_{1}=\sqrt{k_{z}^{2}-\omega^{2}/c^{2}}$ and $\kappa_{2}=\sqrt{k_{z}^{2}-\epsilon_{2}\omega^{2}/c^{2}}$. For a large $a$, the dispersion relation in Eq. (1) resembles for $m=0$ as $k=(\omega/c)\sqrt{\epsilon_{2}/(1+\epsilon_{2})}$, which is the dispersion relation of SPP on the flat surface [14]. In the $m=0$ mode, the phase velocity $v_{ph}$ evolves similarly with that in planar geometry and only approaches $c$ as $n_{e}$ is much higher than $n_{c}$. However, in $m=1$ mode, $v_{ph}$ is very close to $c$ even for a low plasma density. This feature enables the $m=1$ mode to be used for relativistic electron acceleration in a near-critical plasma.
@@
Figure 3: PIC results: 3D (a) and front (b) views of the high-energy ($\mathcal{E}>88\leavevmode\nobreak\ $\mathrm{M}\mathrm{e}\mathrm{V}$$) electron beam modulation after the laser propagates $z=40\leavevmode\nobreak\ $\mu\mathrm{m}$$ in the microtube. The arrows in (b) present the transverse momentum field ($u_{x},u_{y}$). (c): Energy evolution of the high-energy electrons in the beam (grey lines) and the mean value (black line) as a function of propagation distance. (d) Trajectories of these electrons. The red line represents one of the electrons in the beam.
@@
The constant $A$ is proportional to the laser field amplitude $E_{0}$ and the overlap efficiency factor $\mathcal{O}(a,w_{0},\kappa,m=1)$ as $A\approx\mathcal{O}\cdot E_{0}$. From the PIC simulations as shown in Fig. 2, $\mathcal{O}\simeq 0.19$, which implies that $100\leavevmode\nobreak\ $\mathrm{T}\mathrm{W}$$-level laser pulse can drive $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP field. The efficiency is orders of magnitude higher than that with a linearly polarised pulse where $\mathcal{O}\sim 10^{-3}$ [15]. The absolute value of $A$ depends on the incident laser power and the overlap integral between the incident laser field profile at the coupling region and the SPP mode profile, as
@@
where $\perp$ denotes components transverse to the z-axis. Eq. (2) indicates that, except $m=0$ mode, the SPP field with the azimuthal phase dependence $e^{im\phi}$ allows the enhanced coupling if the incident field matches the mode. As a result, the efficient SPP excitation depends on how the laser field matches the azimuthal symmetry of the SPP mode. The coupling efficiency can be determined by $(\mathbf{E}_{spp}^{(m)}\propto e^{im\phi})\cdot(\mathbf{E}_{\text{laser}}^{*}\propto e^{-il\phi})\sim e^{im\phi}\times e^{-il\phi}=e^{i(m-l)\phi}$ where $l$ is the azimuthal index of the laser field. For a CP laser pulse, $l=\pm 1$ and the electric field $\mathbf{E}_{\text{laser}}(r,\phi,z,t)\approx E_{0}(\hat{\mathbf{r}}+i\hat{\boldsymbol{\phi}})g(r)e^{\pm i\phi}e^{i(k_{l}z-\omega_{l}t)}$ where $g(r)$ represents the spatial envelope. The symbol $\pm$ denotes the right (RHC) or left (LHC) hand circular polarisation, respectively. The azimuthal structure of RHC or LHC polarisation can match the azimuthal phase dependence of rotation SPP mode $m=1$ or $m=-1$, as $e^{\pm i\phi}\times e^{\mp i\phi}=1$, respectively. This leads to strongly enhanced coupling of these modes as seen in Fig. 2. For the other modes, e.g. $m\neq 1$, integrating $e^{i(m\pm 1)\phi}$ over $2\pi$ gives zero, and the excitation is not efficient.

### v2

@@ Abstract
Coherent synchrotron radiation (CSR) is crucial for the development of powerful ultrashort light sources. We present a mechanism for generating CSR in the form of generalised superradiance, based on surface plasmon polaritons (SPPs), which are resonantly excited on a solid, near-critical-density inner surface of a microtube. A high-intensity, circularly polarised laser pulse, propagating along the microtube axis, efficiently couples the cylindrical SPP modes. This process creates azimuthally structured, rotating electromagnetic fields. These rotating fields subsequently confine, modulate, and directly accelerate surface electrons to emit CSR in the Vavilov-Cherenkov angle. We further demonstrate that by improving the azimuthal symmetry of these electrons, the helical modulation enables CSR emission across all azimuthal directions in the form of isolated harmonics, significantly enhancing radiation intensity even when full coherence is imperfect. Our full 3D Particle-in-Cell simulations indicate this scheme can generate X-rays with coherence enhanced by up to two orders of magnitude compared to incoherent emission. The challenges to experimentally realise this scheme are discussed, including the need for high-contrast lasers to prevent pre-plasma formation and the demanding tolerances for microtube fabrication and alignment, while these challenges are not beyond the scope of existing or near-future experimental capabilities.
@@
Coherent synchrotron radiation (CSR) represents a unique regime of electromagnetic (EM) radiation emitted when relativistic charged particles—typically electrons—undergo accelerated motion in a collective, spatially ordered configuration [1]. Its exceptional properties, such as high brightness and broadband emission, offer significant advantages for applications requiring high-power radiation sources in the terahertz (THz) and far-infrared regimes, such as ultrafast spectroscopy [2], particle acceleration diagnostics [3, 4], and advanced light sources [5, 6].
@@
The rapid progress in relativistic laser-plasma interactions has promoted the emergence of new compact radiation sources [7, 8, 9, 10]. Furthermore, an unexplored regime of superradiance has been recently discovered where coherent radiation can be obtained by specific modulation of the bunch to create the luminal or superluminal quasiparticle as the radiator [11], which is distinct from the conventional superradiance mechanism [12] and free electron laser (FEL) [13]. A quasiparticle is a collective excitation of a group of underlying particles that behaves as an effective single entity [14], which can travel at any velocity without violating causality. Due to the superluminal motion, the wavefronts of the radiation emitted at different times can intersect to form an optical shock or caustic. At the crossing angle, the spherical waves are phase-aligned and interfere constructively. This new theory unlocks novel coherent radiation in plasma-based accelerators [15].
Figure 1: Schematic: a CP laser pulse (yellow) enters the vacuum channel of a microtube (grey) and excites SPPs while being scattered at the sharp vertical edge. The rotation mode $m=\pm 1$ (green-yellow) can efficiently couple with the laser field and accelerates the trapped electrons(sphere, blue-to-red colour presents the energy) into a spiral motion along the solid surface for the CSR emission (yellow-red cone).
@@
A high-intensity laser pulse interacting with an interface between a metal and a dielectric medium (e.g., air or vacuum) can induce collective oscillations of free electrons. These oscillations, known as surface plasmons (SPs), are characterised by the plasma frequency, $\omega_{p}=\sqrt{4\pi n_{e}e^{2}/m_{e}}$, where $n_{e}$ is electron density, $e$ is electron charge, $m_{e}$ is electron rest mass [16]. The excitation of SPs has been recently proposed for generating TV/m-level wakefields in solid microtubes — a hollow structure composed of nanomaterials (e.g., a longitudinally aligned CNT forest) for relativistic particle acceleration [17]. When these SPs couple with incident EM waves, they form surface plasmon polaritons (SPPs)—hybrid light-electron modes that propagate along the interface, also termed surface plasmon waves [18]. Direct excitation of SPPs on a smooth, infinite interface is inherently inefficient due to the momentum mismatch between the incident light and the SPP modes [19]. However, in finite-sized solid microtargets, sharp edges can serve as effective scatterers, supplying the necessary momentum to bridge this gap and enable efficient SPP excitation [20]. The EM fields associated with SPPs can directly accelerate electrons along the target surface to relativistic energies. Recent studies have shown that irradiating the edge of a flat microtape with a high-intensity, linearly polarised laser pulse can generate strong SPP, producing high-energy electrons and high-flux X-ray radiation [21].
@@
In this letter, we propose a comprehensive, self-contained and highly efficient plasma-based scheme to realise the principle of generalised superradiance, by using SPPs on a cylindrical surface to create a well-controlled field structure with a $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level gradient which accelerates and modulates the self-trapped electron bunch acting as a superluminal quasiparticle. We first show that SPPs can be directly excited on the cylindrical solid interface between plasma and vacuum by a circularly polarised (CP) laser pulse propagating paraxially through a microtube of near-critical density, as illustrated in Fig. 1. The cylindrical geometry enables the CP laser to couple efficiently with a rotating SPP mode whose phase velocity approaches the speed of light in vacuum. This resonant coupling provides $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP fields for the modulated trapping, confinement and stable acceleration of relativistic electrons over several Rayleigh lengths. By matching the channel radius to the SPP field, the modulated, high-energy electrons rotate collectively as a quasiparticle near the surface, which produces CSR at the Vavilov-Cherenkov angle. We further demonstrate that, with an azimuthally symmetric distribution, the modulated beam can emit coherent radiation in all azimuthal directions and allow the harmonics to be well-separated. Full 3D classical particle-in-cell (PIC) simulations have been used to verify the fundamental process. Numerical simulations demonstrate a significant coherence enhancement of photon emission compared to the incoherent case, with a radiation spectrum centred at a critical energy of several $\mathrm{k}\mathrm{e}\mathrm{V}$. These results represent a significant step towards a viable experimental platform from the theoretical concept of generalised superradiance.
@@
Figure 2: (a) Electron density distribution $n_{e}$ (black colour) and $E_{z}$ (blue-red colourmap) in yzyz-plane sliced at $x=0$ while the laser pulse entering the microtube. (b) Snapshots of $E_{z}$ in xyxy-plane sliced at $z=5\penalty 10000\ $\mu\mathrm{m}$\$ corresponding to (a). (c)-(e) Snapshots of $E_{r}$, $E_{\phi}$ and $E_{z}$ in xyxy-plane at the same phase after the laser pulse propagating $z=41\penalty 10000\ $\mu\mathrm{m}$\$ in the tube. The solid black and dashed red lines in each plot present the line plots of the corresponding field along $y=0\penalty 10000\ $\mu\mathrm{m}$\$ from PIC and analytical solutions, respectively.
A microtube can be made of nano-materials, e.g., a longitudinally aligned carbon nanotube (CNT) forest, which can provide the solid plasma with flexible density in the range of $10^{19}-10^{24}\penalty 10000\ $\mathrm{c}\mathrm{m}^{-3}$\$ [22]. The excited SPP and corresponding electric fields are shown in Fig. 2, which are obtained from a PIC simulation where a microtube of radius $a=1.5\penalty 10000\ $\mu\mathrm{m}$\$ is used. A $20\penalty 10000\ $\mu\mathrm{m}$\$-long vacuum section is placed at the head of the target to initialise the laser pulse and provides a sharp annular edge at $z=0\penalty 10000\ $\mu\mathrm{m}$\$. As the bulk size of the tube is much larger than the inner radius of a CNT, e.g. up to several $\mathrm{n}\mathrm{m}$, the effect from the inner structure of CNTs is neglected. A CP laser pulse of wavelength $\lambda_{l}=0.8\penalty 10000\ $\mu\mathrm{m}$\$ and normalised strength $a_{0}=6$ propagates along the axis in the $z$ direction and is focused to the entrance of the microtube, with root-mean-square (RMS) waist $w_{0}=2.0\penalty 10000\ $\mu\mathrm{m}$\$, and duration $\sigma_{\tau}=20.0\penalty 10000\ $\mathrm{f}\mathrm{s}$\$. These parameters give a peak field of $24.0\penalty 10000\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$\$, or $13.6\penalty 10000\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$\$ on the inner surface. This is sufficient for the pulse front to instantaneously ionise electrons of the material within half a cycle. Therefore, the material properties, e.g., lattice structure or optical absorption, are unimportant for electron dynamics. For femtosecond laser pulses, the interaction time is also in the order of femtoseconds. Significant hydrodynamic surface expansion, however, occurs on a much slower picosecond timescale. Therefore, for moderate laser strength with high contrast, the solid surface can be maintained during the interaction. These conditions allow the propagation of SPPs on the vacuum-plasma interface.
The simulations are carried out with the code WarpX [23]. The dimensions of the moving window are $8\penalty 10000\ $\mu\mathrm{m}$\times 8\penalty 10000\ $\mu\mathrm{m}$\times 30\penalty 10000\ $\mu\mathrm{m}$\$, comprising $512\times 512\times 1024$ cells in the $x$, $y$, and $z$ directions, respectively. Each cell contains 64 macro particles, which are sufficient to solve the electron dynamics effectively. The microtube is initially modelled as a uniform distribution of neutral carbon atoms, with their ionisation energies adjusted to reflect CNT properties—specifically the weak $\pi$-bond and $\sigma$ (C–C) bond energies. The initial carbon density is $n_{C0}=2\times 10^{21}\penalty 10000\ $\mathrm{c}\mathrm{m}^{-3}$\$ which gives $\omega_{p}$ close to the laser frequency $\omega_{l}$. The field ionisation is implemented using the Ammosov-Delone-Krainov (ADK) method [24].
@@
The vertical solid surface at $z=0\penalty 10000\ $\mu\mathrm{m}$\$ scatters the incident CP laser pulse and enables the direct matching with the SPP mode if the resonant condition $k_{l}=k_{spp}$ is satisfied, where $k_{l}$ and $k_{spp}$ are wave numbers of laser and SPP field, respectively. This condition is efficient near critical density, as shown in Fig. 2(a) and (b). The excited SPP fields can stably propagate with ultrahigh strength in $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level, e.g. $4.5\penalty 10000\ $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$\$ in our simulations as shown in Fig. 2 (c)-(e). Some electrons can be trapped into the SPP field and accelerated to relativistic speed by the longitudinal component $E_{z}$ and, at the same time, confined in a small vicinity near the surface by the transverse component $E_{r}$. The azimuthal component $E_{\phi}$ drives the modulated beam to rotate.
@@
The PIC results can be confirmed theoretically by analysing the SPP excitation in cylindrical geometry, $(r,\phi,z)$ [25]. In general, a pure transverse magnetic (TM) mode (e.g. $B_{spp,z}=0$ and $E_{spp,z}\neq 0$) of SPP field can exist on the cylindrical surface. The $z$-component of a SPP mode can be given by solving Maxwell equations, as $E_{spp,z}^{(m)}(r,\phi,z)=Af_{m}(\kappa r)e^{i(k_{z}z+m\phi-\omega t)}$, where $A$ is a constant representing the field amplitude. $m$ is the SPP mode number and $\kappa=\sqrt{k_{z}^{2}-\epsilon_{d}\omega^{2}/c^{2}}$ with $k_{z}$ the wavevector along the tube axis, $\omega$ the frequency, $c$ is the speed of light in vacuum and $\epsilon_{d}$ permittivity of the medium. $f_{m}(\kappa r)$ presents the radial field profile and depends on the medium properties. In the vacuum channel, $r<a$, the field is presented by the modified Bessel functions of the first kind as $f_{m}(\kappa r)=I_{m}(\kappa_{1}r)$ since the field is finite as $r\to 0$ and $\epsilon_{d}=\epsilon_{1}=1$. In the plasma, $r>a$, the field is presented by the the modified Bessel functions of the second kind $f_{m}(\kappa r)=K_{m}(\kappa r)$ as the field decays to zero as $r\to\infty$ and $\epsilon_{d}=\epsilon_{2}=1-n_{e}/n_{c}(\omega)$ where $n_{c}$ is the critical density for the frequency $\omega$ as $n_{c}(\omega)=m_{e}\omega^{2}/4\pi e^{2}$. PIC simulation agrees well with the analytical solution as shown in Fig. 2(c)-(e). The other components can be obtained from the axial components using Maxwell’s equations, for example, the electric components as $E_{spp,r}^{(m)}=(ik_{z}/\kappa^{2})\partial E_{spp,z}^{(m)}/\partial r$ and $E_{spp,\phi}^{(m)}=imk_{z}E_{spp,z}^{(m)}/\kappa^{2}r$ for different regions, respectively. The dispersion relation can be derived by considering the continuity of the SPP field at the interface of plasma and vacuum. It shows that in $m=1$ mode, $v_{ph}$ is very close to $c$ even for a low plasma density. This feature enables the $m=1$ mode to be used for relativistic electron acceleration in a near-critical plasma. See more details in Appendix A and the Supplemental Material [26, 27, 28].
@@
Figure 3: PIC results: 3D (a) and front (b) views of the high-energy ($\mathcal{E}>88\penalty 10000\ $\mathrm{M}\mathrm{e}\mathrm{V}$\$) electron beam modulation after the laser propagates $z=40\penalty 10000\ $\mu\mathrm{m}$\$ in the microtube. The arrows in (b) present the transverse momentum field ($u_{x},u_{y}$). (c): Energy evolution of the high-energy electrons in the beam (grey lines) and the mean value (black line) as a function of propagation distance. (d) Trajectories of these electrons. The red line represents one of the electrons in the beam.
@@
The constant $A$ is proportional to the laser field amplitude $E_{0}$ and the overlap efficiency factor $\mathcal{O}(a,w_{0},\kappa,m=1)$ as $A\approx\mathcal{O}\cdot E_{0}$. From the PIC simulations as shown in Fig. 2, $\mathcal{O}\simeq 0.19$, which implies that $100\penalty 10000\ $\mathrm{T}\mathrm{W}$\$-level laser pulse can drive $\mathrm{T}\mathrm{V}\mathrm{/}\mathrm{m}$-level SPP field. The efficiency is orders of magnitude higher than that with a linearly polarised pulse, where $\mathcal{O}\sim 10^{-3}$ [17]. The absolute value of $A$ depends on the incident laser power and the overlap integral between the incident laser field profile at the coupling region and the SPP mode profile, as
@@
where $\perp$ denotes components transverse to the z-axis. Eq. (1) indicates that an efficient overlap requires $w_{0}$ to be comparable with the channel size $a$. This enables a larger laser focus to be matched with a wider tube. It also indicates that, except $m=0$ mode, the SPP field with the azimuthal phase dependence $e^{im\phi}$ allows the enhanced coupling if the incident field can match the mode. As a result, the efficient SPP excitation depends on how the laser field matches the azimuthal symmetry of the SPP mode. The coupling efficiency can be determined by $(\mathbf{E}_{spp}^{(m)}\propto e^{im\phi})\cdot(\mathbf{E}_{\text{laser}}^{*}\propto e^{-il\phi})\sim e^{im\phi}\times e^{-il\phi}=e^{i(m-l)\phi}$ where $l$ is the azimuthal index of the laser field. For a CP laser pulse, $l=\pm 1$ and the electric field $\mathbf{E}_{\text{laser}}(r,\phi,z,t)\approx E_{0}(\hat{\mathbf{r}}+i\hat{\boldsymbol{\phi}})g(r)e^{\pm i\phi}e^{i(k_{l}z-\omega_{l}t)}$ where $g(r)$ represents the spatial envelope. The symbol $\pm$ denotes the right (RHC) or left (LHC) hand circular polarisation, respectively. The azimuthal structure of RHC or LHC polarisation can match the azimuthal phase dependence of the rotation SPP mode $m=1$ or m=1m=-1, as $e^{\pm i\phi}\times e^{\mp i\phi}=1$, respectively. This leads to strongly enhanced coupling of these modes as seen in Fig. 2. For the other modes, e.g. $m\neq 1$, integrating $e^{i(m\pm 1)\phi}$ over $2\pi$ gives zero, and the excitation is not efficient. This edge-coupling mechanism presented in this letter also enables the excitation of other modes by matching the azimuthal mode number $m$, for example, by a Laguerre-Gaussian laser [29].